Can you solve each system by addition.
x+y=7
x-y=9

Answer by unlockmath(1122)

Hello,
Let's add both equations.
x+y=7
x-y=9 It will look like this:
2x=16 (Notice the y cancels out)Now divide both sides by 2
x=8 Plug 8 into the original equation (either one)
8+y=7 Subtract 8 from both sides gives us:
y=-1 There we go. You can check these by plugging in the answers into the equations.
RJ Toftness
www.math-unlock.com